I'm having trouble with this problem. If anyone can help and explain how to get the answer that would be greatly appreciated. Have to take a test on thurs. "Point A(-4,1) is in the standard (x,y) coordinate plane. What must be the coordinates of point B so that the line x=2 is the perpendicular bisector of AB?" Thanking you in advance.
Answer by stanbon(75887)

Point A(-4,1) is in the standard (x,y) coordinate plane.
What must be the coordinates of point B so that the line x=2 is the perpendicular bisector of AB?"
-----
Plot the point (-4,1)
Draw the line x = 2
---------------------------
Notice that the line passes thru (2,1)
That point is the midpoint of AB
---
Let (x,y) be the coordinates of "B".
--------------------------------------
Solve for x and for y:
(x+(-4))/2 = 2
x-4 = 4
x = 8
---------
(y+1)/2 = 1
y+1 = 2
y = 1
----------------------
The coordinates of "B" are(8,1)
======================================
Cheers,
Stan H.